OpenStudy (anonymous):

condense 4 ln x - 5 ln x Please show work. Thanks!

OpenStudy (mertsj):

\[\ln\frac{a}{b}=\ln(a)-\ln(b)\]

OpenStudy (mertsj):

And \[\ln(a)^n=nln(a)\]

OpenStudy (mertsj):

so \[4\ln(x)-5\ln(x)=\ln\frac{x^4}{x^5}=\ln(x)^{-1}\]